# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ Abstract
While Graph Neural Networks (GNNs) have proven highly effective at modeling relational data, pairwise connections cannot fully capture multi-way relationships naturally present in complex real-world systems. In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations –such as simplicial or cellular complexes– to accommodate higher-order interactions. Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism. To overcome these limitations, we propose HOPSE (Higher-Order Positional and Structural Encoder)—a message passing-free framework that uses Hasse graph decompositions to derive efficient and expressive encodings over arbitrary higher-order domains. Notably, HOPSE scales linearly with dataset size while preserving expressive power and permutation equivariance. Experiments on molecular, expressivity and topological benchmarks show that HOPSE matches or surpasses state-of-the-art performance while achieving up to 7× speedups over HOMP-based models, opening a new path for scalable TDL.
@@ 1 Introduction
Graph neural networks (GNNs) [46, 11] have demonstrated impressive capabilities in modeling relational data structured as graphs. However, their reliance on edges to encode interactions inherently limits them to pairwise relationships, which may not fully capture the complexity of many real-world systems. For instance, social dynamics often involve group interactions, while physical phenomena like electrostatic interactions in proteins can span multiple atoms simultaneously.
@@
To address such limitations, Topological Deep Learning (TDL) [8, 23] introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes. Within TDL, standard Topological Neural Networks (TNNs) [18, 12, 22, 21, 38] extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation. [^1] The terms higher-order topological domain and combinatorial representation are used intercheangeably: please see Section A.1 for the formal definitions. This generalization has shown strong performance across diverse fields—including social networks [30], protein biology [28], physics-informed learning [45], and computer networks [6, 5]—where interactions naturally extend beyond pairwise connections.
@@
However, despite their theoretical advantages, TNNs face substantial challenges stemming from the HOMP paradigm. Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain –where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected– leads to an explosion in the number of possible sequences of message passing routes across these different levels [42, 33, 37]. More specifically, for a combinatorial complex $\mathcal{T}$ with dimension –i.e. maximum rank– $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix B for a detailed analysis).
@@
On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs [37, 20]. This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability, posing a major barrier to their deployment in large-scale or time-sensitive real-world applications.
In response to these challenges, this work introduces Higher-Order Positional and Structural Encoder (HOPSE), a novel framework designed to overcome the computational hurdles associated with HOMP. Specifically, HOPSE main features and contributions include:
  • Higher-order topological encodings: By exploiting Hasse graph decompositions –i.e. transformations of combinatorial representations into simpler graph ensembles [23, 38]–, HOPSE facilitates the efficient use of established graph positional and structural encoding techniques (e.g. [13]) to obtain analog representations for higher-order topological domains.
  • General message passing-free method: Inspired by SaNN model [20], the proposed encoding-based methodology effectively captures relational information without explicit message passing mechanisms. Moreover, HOPSE is no longer constrained to simplicial complexes; by design, it can be applied to any higher-order combinatorial representations.
@@
  • Scalability and computational efficiency: In contrast to HOMP-based networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section 4.3). This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section 5 for experimental details).
  • Theoretical guarantees: Even without message passing, HOPSE is as expressive as existing state-of-the-art HOMP-based models (e.g. Papillon et al. [38]) in distinguishing non-isomorphic complexes (Theorem 4.1), while preserving cell permutation equivariance (Theorem 4.2).
  • Flexible instantiations: Two HOPSE realizations are presented—HOPSE-M, relying on handcrafted graph-theoretical encodings, and HOPSE-G, using a pretrained GPSE encoder—offering complementary trade-offs between preprocessing cost and representational richness (Section 4.3).
  • Extensive empirical validation: Comprehensive experiments on molecular benchmarks [34], CLS expressivity dataset [36], and MANTRA [2] –recent dataset consisting of pure topological tasks– validate the practical effectiveness of HOPSE, showcasing its ability to efficiently capture complex relational structures and outperform existing baselines –while being up to 7×7$\times$7 × faster than HOMP-based models (Section 5).
@@
Additionally, Appendix B provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
@@ 2 Background
This section introduces key concepts underpinning TDL. For brevity and generality, the following discussion centers on combinatorial complexes, which subsume all of the discrete topological domains leveraged by TDL [23]. Appendix A provides rigorous mathematical definitions of these other domains, such as simplicial and cellular complexes, along with an extended theoretical background.
@@
The elements of $\mathcal{V}$ represent the nodes, while the elements of $\mathcal{C}$ are called cells (i.e., groups of nodes). The cell $\sigma\in\mathcal{C}$ of rank $k:=\textrm{rk}(\sigma)$ is referred to as a $k$-cell. $\mathcal{C}$ simplifies notation for $(\mathcal{V},\mathcal{C},\textrm{rk})$. Lastly, the dimension of ${\mathcal{T}}$ is defined as the maximal rank among its cells: $\mathrm{dim}(\mathcal{C}):=\max_{\sigma\in\mathcal{C}}\textrm{rk}(\sigma)$.
Signals over combinatorial complexes.
A combinatorial complex $\mathcal{T}=(\mathcal{V},\mathcal{C},\mathrm{rk})$ can be equipped with signals, generalizing the concept of featured graph [^2] Attributed graph and featured graph are used interchangeably to comply with the literature. (see Definition A.1) to higher-order structures. In this context, a signal is a set of functions =$\mathcal{F}=\{F_{r}\}_{r\geq 0}$, where $F_{r}:\{\sigma\in\mathcal{C}\mid\mathrm{rk}(\sigma)=r\}\to\mathbb{R}^{D_{r}}$, where $D_{r}$ is the dimension of the feature space for rank-$r$ cells. For this work, it is important to emphasize that a complex can have multiple attribute functions over each $r$-cell. More formally, =$\mathcal{F}=\{F_{r,k}\}_{k=1}^{K}$ for all $r\geq 0$, where each $F{r,k}$ shares the same domain: $\text{Dom}(F_{r,k})=\text{Dom}(F_{r})$.
@@
Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces rich topological structures. Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex ${\mathcal{T}}$ maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$. Traditional neighborhood functions include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ –through adjacency relations w.r.t. cells of another rank $s$–, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks –from a source rank $s$ to a target rank $t$. Figure 1 presents some instances, and Section A.2 provides formal definitions.
Figure 1: Neighborhood expansions of a complex. Given a complex ${\mathcal{T}}$ (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhoods: adjacency of nodes w.r.t. edges (${\mathcal{A}}_{0,1}$), incidence from nodes to edges (${\mathcal{I}}_{0\to 1}$), incidence from faces to nodes (${\mathcal{I}}_{2\to 1}$), and incidence from nodes to faces (${\mathcal{I}}_{0\to 2}$).
@@
Given a complex ${\mathcal{T}}$, each particular neighborhood $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ [39], defined as the directed graph whose nodes and edges are given, respectively, by
@@
Figure 1 visually shows some examples of these neighborhood graph expansions. Notably, “strictly” comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to ${\mathcal{N}}$. In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ [23].
@@ 3 Related works
Positional and Structural Encodings in GNNs.
@@
In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations. Positional encodings typically capture a node’s location within the graph (e.g., via distances or coordinates), while structural encodings describe the node’s role or function in the topology (e.g., node degrees). Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) [17], while a common structural encoding is the random-walk embedding (RWSE) [15, 43], which reflects multi-hop connectivity; these handcrafted features help differentiate nodes with otherwise similar local neighborhoods. More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms [14, 40, 31]. Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually [14]. Building upon this, the Graph Positional and Structural Encoder (GPSE) [13] offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining. GPSE’s embeddings can be integrated into downstream GNNs, achieving performance on par with or exceeding task-specific encodings, and advancing the development of general-purpose graph models.
Combinatorial Complex Neural Networks.
To overcome the pairwise limitations of graphs, recent work has turned to explicit representation of higher-order interactions through topological domains. Combinatorial complexes provide a unified framework for modeling higher-order interactions with hierarchical organization, and their associated Combinatorial Complex Neural Networks (CCNNs) enable message passing across cells of varying rank (HOMP paradigm), naturally capturing multi-way relationships [23]. Generalizing and subsuming CCNNs, Generalized CCNNs (GCCNs) [39] offer a more flexible and powerful framework: they systematically decompose a complex into a set of strictly augmented Hasse graphs –each one capturing a particular neighborhood– in a modular way that preserves the higher-order structure. In doing so, GCCNs not only facilitate the design of new TNN architectures but also help bridge the gap between graph deep learning and TDL. However, existing TNNs –CCNNs and GCCNs alike– suffer from the high computational complexity associated to HOMP, limiting their practical use in large real-world applications.
Positional encodings in higher-order domains.
Encoding positional information for higher-order topological elements poses unique challenges. Unlike in graphs, where positional features are typically assigned to nodes, each cell rank of a combinatorial complex –nodes, edges, faces,…– potentially requires its own descriptor. Recent works have begun addressing this, though this topic still remains largely unexplored. For instance, the Cellular Transformer (CT) [1] extends transformer architectures to cell complexes by incorporating attention mechanisms that respect higher-order incidence relations. To support this, the authors introduce a set of manually crafted positional encodings tailored to cellular complexes, though their results did not shed light on how to build robust encodings—different ones perform best on different tasks. Moreover, CT incurs quadratic complexity in the number of cells due to the attention mechanism, limiting its scalability to medium-to-large scenarios. In contrast, Simplicial-aware Neural Networks (SaNN[20] take a more efficient route by bypassing message passing entirely. SaNN precomputes structural and positional summaries (e.g., over lower, upper, and co-neighbors) and feeds them as fixed-length vectors into a standard neural network. This achieves constant runtime and memory overhead while maintaining strong expressivity and performance. Nevertheless, the SaNN architecture is also constrained in terms of the topological domain (simplicial complex-specific), as well as with the choices of neighborhoods –which haves an impact into its theoretical expressivity and applicability.
@@ 4 HOPSE: higher-order positional and structural encoder
Section 4.1 outlines the proposed HOPSE pipeline, Section 4.2 establishes its theoretical guarantees, and Section 4.3 presents two practical realizations of the framework.
Let ${\mathcal{T}}$ be an arbitrary combinatorial complex, and ${\mathcal{N}}_{C}\subset{\mathcal{N}}_{\mathcal{T}}$ any subset of neighborhoods on ${\mathcal{T}}$. The following notation is introduced to support the upcoming presentation of the framework:
Strictly augmented multi-attributed Hasse graphs.
For a given neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{\mathcal{C}}$, its strictly augmented multi-attributed Hasse graph –denoted by ${\mathcal{J}}_{{\mathcal{N}}}$– assigns to the resulting strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ the corresponding signals $\mathcal{F}_{{\mathcal{N}}}$ over the cells present in that neighborhood, i.e. ${\mathcal{J}}_{{\mathcal{N}}}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},\mathcal{F}_{\mathcal{N}})$. The shorthand ${\mathcal{F}}_{{\mathcal{N}},k}=\{F_{r,k}\}_{r\geq 0}$ is the set of all $k$-attribute functions for each rank $r$ that act on the subset of cells in a neighborhood ${\mathcal{C}}_{\mathcal{N}}$ as $F_{r,k}:\{\sigma\in{\mathcal{C}}_{\mathcal{N}}\}\to\mathbb{R}^{D_{r}}$. To streamline the management of multiple associated signals, the restriction of ${\mathcal{J}}_{{\mathcal{N}}}$ to the $k$-th feature is defined as $\evaluated{{\mathcal{J}}_{{\mathcal{N}}}}_{k}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},{\mathcal{F}}_{{\mathcal{N}},k})$. Lastly, $\mathbf{J}_{C}=\left\{{\mathcal{J}}_{{\mathcal{N}}}\right\}_{{\mathcal{N}}\in{\mathcal{N}}_{C}}$ denotes the corresponding Hasse graph decomposition of ${\mathcal{T}}$ w.r.t. a given neighborhood set ${\mathcal{N}}_{C}$.
Rank-targeted neighborhoods.
The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.

### v2

@@ Abstract
While Graph Neural Networks (GNNs) have proven highly effective at modeling relational data, pairwise connections cannot fully capture multi-way relationships naturally present in complex real-world systems. In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations —such as simplicial or cellular complexes— to accommodate higher-order interactions. Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism.This work presents HOPSE (Higher-Order Positional and Structural Encoder), an alternative method to solve tasks involving higher-order interactions without message passing. Instead, HOPSE breaks apart arbitrary higher-order domains into it’s neighborhood relationships using a Hasse graph decompositions. This method shows that decoupling the representation learning of the topology of each neighborhood and the attributes results in a lower computational complexity, putting into question the need for HOMP.The experiments on molecular graph tasks and topological benchmarks show that HOPSE matches performance in traditional TDL datasets and outperforms HOMP methods in topological tasks, all while achieving up to $7\times$ speedups over HOMP-based models, opening a new path for scalable TDL.
@@ 1 Introduction
Graph neural networks (GNNs) (Zhou et al., 2020; Bronstein et al., 2021) have demonstrated impressive capabilities in modeling relational data structured as graphs. However, their reliance on edges to encode interactions inherently limits them to pairwise relationships, which may not fully capture the complexity of many real-world systems. For instance, social dynamics often involve group interactions, while physical phenomena like electrostatic interactions in proteins can span multiple atoms simultaneously.
@@
To address such limitations, Topological Deep Learning (TDL) (Bodnar, 2023; Hajij et al., 2023) introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes. Within TDL, standard Topological Neural Networks (TNNs) (Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; 2022; Papillon et al., 2023) extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation. [^1] The terms higher-order topological domain and combinatorial representation are used intercheangeably: please see Section A.1 for the formal definitions. This generalization has shown strong performance across diverse fields including social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics-informed learning (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025; 2023) where interactions naturally extend beyond pairwise connections.
@@
However, despite their theoretical advantages, TNNs face substantial challenges stemming from the HOMP paradigm. Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain, where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected, leads to an explosion in the number of possible sequences of message passing routes across these different levels (Telyatnikov et al., 2024; Montagna et al., 2024; Papamarkou et al., 2024). More specifically, for a combinatorial complex $\mathcal{T}$ with dimension —i.e. maximum rank— $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix˜C for a detailed analysis).
@@
On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs (Papamarkou et al., 2024; Gurugubelli and Chepuri, 2024). This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability and is not required for many tasks, even ones that require leveraging higher-order relationship. Appendix˜B contains theoretical motivation for this claim.
In response to these challenges, this work introduces the Higher-Order Positional and Structural Encoder (HOPSE), a novel and scalable framework that bypasses the HOMP paradigm altogether. While the message-passing preprocessing philosophy has been successfully applied within the simplicial domain (SaNN (Gurugubelli and Chepuri, 2024)), HOPSE does not rely on it altogether and achieves a critical generalization to arbitrary higher-order domains. By employing the Hasse graph decomposition as a methodological bridge, this framework enables the rich literature of powerful graph-based positional and structural encodings (PSEs) to be systematically and flexibly applied to any higher-order structure.
More specifically, HOPSE’s main features and contributions include:
  • Higher-order topological characterization: By decomposing a combinatorial complex into simpler graph ensembles, called Hasse graphs  (Hajij et al., 2023; Papillon et al., 2023), HOPSE efficiently uses established graph PSE techniques (e.g. (Cantürk et al., 2023)) to obtain characterizations for higher-order topological domains.
@@
  • Scalability and computational efficiency: In contrast to HOMP networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section˜4.2). This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section˜5 for experimental details).
  • Extensive empirical validation: Comprehensive experiments on graph  (Morris et al., 2020) and topological  (Ballester et al., 2025) datasets validate the practical effectiveness of HOPSE. It is able to perform on-par to HOMP methods on graph tasks while showing stronger characterization capabilities required for topological tasks with a speed up of up to $7\times$ with respect to HOMP models (Section˜5).
@@
Additionally, Appendix˜C provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
@@ 2 Background
This section introduces key concepts underpinning TDL. For brevity and generality, the following discussion centers on combinatorial complexes (CCs), which subsume all of the discrete topological domains leveraged by TDL Hajij et al. (2023). Appendix A provides rigorous mathematical definitions of these other domains, such as simplicial and cellular complexes, along with an extended theoretical background.
@@
The elements of $\mathcal{V}$ represent the nodes, while the elements of $\mathcal{C}$ are called cells (i.e., groups of nodes). The cell $\sigma\in\mathcal{C}$ of rank $k:=\textrm{rk}(\sigma)$ is referred to as a $k$-cell. Lastly, the dimension of $\mathcal{T}$} is defined as the maximal rank among its cells: $\mathrm{dim}(\mathcal{C}):=\max_{\sigma\in\mathcal{C}}\textrm{rk}(\sigma)$.
@@
Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces topological structures. Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex $\mathcal{T}$} maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}$. Traditional neighborhood functionsme include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ —through adjacency relations w.r.t. cells of another rank ss—, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks from a source rank $s$ to a target rank $t$. Figure˜1 presents some instances, and Section˜A.2 provides formal definitions.
Figure 1: Neighborhood expansions of a complex. Given a complex $\mathcal{T}$} (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhood functions: adjacency of nodes w.r.t. edges (${\mathcal{A}}_{0,1}$), incidence from nodes to edges (${\mathcal{I}}_{0\to 1}$), incidence from faces to nodes (${\mathcal{I}}_{2\to 1}$), and incidence from nodes to faces (${\mathcal{I}}_{0\to 2}$).
@@
Given a complex $\mathcal{T}$}, each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ (Papillon et al., 2024), defined as the directed graph whose nodes and edges are given, respectively, by
@@
Figure˜1 visually shows some examples of the graph expansions induced by these neighborhood functions. Notably, strictly comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to $\mathcal{N}$}. In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ (Hajij et al., 2023).
Rank-targeted neighborhood functions.
The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{S}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.
@@ 3 Related works
Graph characterization in GNNs.
@@
In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations and characterize graphs. Positional encodings typically capture a node’s location within the graph (e.g., via distances or coordinates), while structural encodings describe the node’s role or function in the topology (e.g., node degrees). Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) Dwivedi et al. (2023), while a common structural encoding is the random-walk embedding (RWSE) Dwivedi et al. (2022); Wang et al. (2022), which reflects multi-hop connectivity. These handcrafted features help characterize graphs with otherwise similar local neighborhoods. More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms Dwivedi et al. (2021); Rampášek et al. (2022); Ma et al. (2023). Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually (Dwivedi et al., 2021). Building upon this, the Graph Positional and Structural Encoder (GPSE) Cantürk et al. (2023) offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining. GPSE’s embeddings can be integrated into downstream GNNs, achieving performance on par with or exceeding task-specific encodings, and advancing the development of general-purpose graph models.
Efficient message passing methods
There are many different ways to achieve expressive and computationally efficient GNNs but the same cannot be said for TNNs. Various works in the GNN domain have approached message passing over subgraphs or partitions of the original graph (Zeng et al., 2021; Li et al., 2021). Additionally, other approaches focus on decoupling learning the features from the structure (Frasca et al., 2020; Doshi and Chepuri, 2022; Chen et al., 2020). In higher-order domains partitions are not trivial, so other methods have been applied to achieve efficiency.Simplicial Scattering Networks  (Madhu et al., 2024), leverage random walks over the intitial simplicial complex which give node structure-informed embeddings resulting in a parameter-free network. However, their experimental framework is mostly focused in node-level tasks. To the best of our knowledge Simplicial-aware Neural Networks (SaNN(Gurugubelli and Chepuri, 2024) is the only method that proposes efficient computations for higher-order networks in a learning scenario. They precompute a deterministic transformation over both the structure and the features of neighborhoods in a simplicial complex. Effectively, this results in a precomputation of message passing. However, it has several drawbacks. First, there is no clear indication of which transformation of the features and structure of a neighborhood should be used. Since this is a preprocessing step, this function cannot be learned. Second, this approach is limited to only simplicial complexes. HOPSE joins the ideas from (Frasca et al., 2020; Chen et al., 2020) and (Gurugubelli and Chepuri, 2024) to partially decouple learning the structure from the attributes. Additionally, HOPSE bypasses the theoretical limitations of (Gurugubelli and Chepuri, 2024) (and HOMP-based models), which is at most as expressive as SWL.
@@ 4 HOPSE: higher-order positional and structural encoder
Section˜4.1 outlines the proposed HOPSE pipelineand Section˜4.2 presents two practical realizations of the framework.